0.00 points A lelephone companys go is 10 have no more i an 4 men i re la ures on any 10 mi es ๙ ine. The compar o me ir e e er ces an average o 6 monui i e aiures per 0 miles o ne Let x de 0 e lhe n moer o mon í a) Find the probabllty thar the company wil moet its goal on a particular 100 miles of Inc (Do not round Intenmedlate calculations.Round Inal answer to 4 decimal places. Probabllty Ib} Find thたpmboblity that thf:cnmpany wil not mnnt its gnal r:n ה parrillar lan mila5of Ire {Do not round intermadiate calculations. Round fira1answer to 4 decimal pacas.) Probabllts c) Fi trve prctiabalty ha te campary wil tave no mare than 4 mvanthly taurs con a prarticutar 200 mis ine.(Do not round intermediate calculations. Round tinal answer to 4 deoimal places. re iues per 100 miles o ne ss mi x r as a Poisson (d} I rd ute prubetlitr Ural utecorrally enl hie rnule utlan 12 munu ly i비ures un aluculai 15ข mies ol ine. Do not round intermedite calculations. Round final answer to 4 decimal places.) References eBook & Resources Delliculy. 3 Haid Setting Windows
Answer #1

Solution:

(a) Find the probability that the company will meet the goal on a particular 100 miles of line?

Answer: Let X denotes the number of monthly line failures per 100 miles of line. Then we have:

λ--x 100-12 50

Therefore, we have to find here:

P(X < 4

Using the Poisson distribution, we have:

P(X < 4) = 0.0076

. Probability 0.0076

(b) Find the probability that the company will not meet the goal on a particular 100 miles of line?

Answer: Probability P(X > 4) = 1-P(X < 4) 1-0.0076= 0.9924

(c) Find the probability that the company will have no more than 4 monthly failures on a particular 200 miles of line?

Answer: Let X denotes the number of monthly line failures per 200 miles of line. Then we have:

λ--x 200-24 50

Therefore, we have to find here:

P(X < 4

Using the Poisson distribution, we have:

P(X < 4) 0.0000

Probability 0.0000

(d) Find the probability that the company will have more than 12 monthly failures on a particular 150 miles of line?

Answer: Let X denotes the number of monthly line failures per 150 miles of line. Then we have:

λ х 150 18 50

Therefore, we have to find here:

P(X> 12)

Using the Poisson distribution, we have:

P(X > 12) = 1-P(X < 12) 1-0.0917-0.9083

small oldsymbol{ herefore Probability =0.9083}